3X-5Y=-9
5X-6Y=-8

Answer by Jk22(389)

Solve :

3X-5Y=-9 | times 5
5X-6Y=-8 | times 3, sub
---------------
-25Y-(-6)*3Y=-45+24
-25Y+18Y=-21
-7Y=-21
---------------
Y=3
X=(-9+15)/3=2
----------------
Verif :

3*2-5*3=6-15=-9.
5*2-6*3=10-19=-8.